# Disentangling, Amplifying, and Debiasing: Learning Disentangled Representations for Fair Graph Neural Networks

### v1

@@ 4 Evaluation
Datasets. We used 5 real-world graph datasets for our experiments: NBA, Recidivism, Credit, Pokec_n, and Pokec_z, which are all publicly available. [^4] https://github.com/chirag126, https://github.com/EnyanDai Table 1 provides key statistics for these datasets (more details in Section 2).
@@
Evaluation Tasks. Following previous studies (Agarwal, Lakkaraju, and Zitnik 2021; Dai and Wang 2021; Dong et al. 2022; Wang et al. 2022), we assessed the methods by using a node classification task, splitting the nodes into training (50%), validation (25%), and test (25%) sets. We measured accuracy with three metrics: Accuracy (ACC), Area Under the Curve (AUC), and F1-Score. In addition, we evaluate fairness by using Statistical Parity (SP) (Dwork et al. 2012) and Equality of Opportunity (EO) (Hardt, Price, and Srebro 2016), where lower values indicate better model fairness. Detailed equations for calculating SP and EO are provided in Section A of the supplementary material.
Implementation Details. We used Vanilla GCN as the backbone for both the competitors and DAB-GNN, carefully tuning their hyperparameters via grid search. For DAB-GNN, we used a 3-layer GCN with a hidden layer of size 16, a gradient penalty (Gulrajani et al. 2017) of 10, 1,000 epochs, the embedding dimensionality of 48, and a weight decay of 0.00001. The hyperparameters $k$ for $k$-NN graph construction and $\alpha$ for $\mathcal{L}_{\text{fh}}$ were tuned within the range of {1, 100}, while $\beta$ for $\mathcal{L}_{\text{bco}}$ was tuned within {0.00001, 0.1}. All experiments were conducted by using five different seed settings, and we report the average accuracy. For complete implementation details, please refer to Section B of the supplementary material.
@@ Results
Figure 2: Visualization of disentangled node embeddings by using t-SNE: AbEmb (purple), SbEmb (red), and PbEmb (yellow). For additional results on other datasets, see Section C of the supplementary material.
(a) Recidivism (b) Credit
@@
Figure 3: The effects of $\alpha$ and $\beta$ on AUC ($\uparrow$) and EO ($\downarrow$). For additional results on other datasets, see Section C of the supplementary material.
@@ 5 Conclusion
In this work, we identified a significant challenge inherent to existing fairness-aware GNN methods: the entanglement of different bias types in the final node embeddings leads to difficulty in their comprehensive debiasing. To address this challenge, we introduced DAB-GNN, a novel GNN framework that disentangles, amplifies, and debiases the attribute, structure, and potential biases within node embeddings. Extensive experiments on five real-world graph datasets demonstrated that DAB-GNN consistently outperforms ten state-of-the-art competitors in balancing accuracy and fairness, while also validating the effectiveness of our design choices in delivering fair and accurate outcomes.
Appendix A Detailed Equations
Fairness Analysis Metrics
    $b_{attr}=\frac{1}{M}\sum_{m=1}^{M}W(\text{pdf}(X_{m}^{0}),\text{pdf}(X_{m}^{1})),$
    where $X_{m}^{i}$ represents the values of the $m$-th dimension (1 $\leq$ $m$leq $M$) of $X$ for nodes in subgroup $i$, and $W$ denotes the Wasserstein-1 distance between the probability density functions (PDFs) of two subgroups.
    $b_{stru}=\frac{1}{M}\sum_{m=1}^{M}W(\text{pdf}(R_{m}^{0}),\text{pdf}(R_{m}^{1})),$
    where $\mathbf{R}=\mathbf{M}_{H}\mathbf{X}_{norm}$ represents the reachability matrix. Here, $\mathcal{R}^{0}_{m}$ and $\mathcal{R}^{1}_{m}$ correspond to the $m$-th attribute value sets in $\mathbf{R}$ for nodes with sensitive attributes of 0 and 1, respectively. To compute $\mathbf{R}$, we define a normalized adjacency matrix with re-weighted self-loops as $\mathbf{P}_{norm}=\alpha\mathbf{A}_{norm}+(1-\alpha)\mathbf{I}$, where $\mathbf{A}_{norm}$ and $\mathbf{I}$ represent the symmetric normalized adjacency matrix and the identity matrix, respectively; $\alpha$ is a hyperparameter ranging from 0 to 1. The propagation matrix is given by $\mathbf{M}_{H}=\sum_{h=1}^{H}\beta_{h}\mathbf{P}_{norm}^{h}$, where $H$ indicates the number of hops and $\beta$ indicates a discount factor that reduces the weight of propagation from more distant neighbors.
    $h_{G}=\frac{\sum_{i}C_{ii}}{\sum_{i,j}C_{ij}},$
    where $C_{ij}$ represents the number of edges between nodes in subgroups $i$ and $j$. A value close to 1 indicates a high concentration of intra-subgroup edges, while a value near 0 indicates more inter-subgroup edges.
    $F_{G}=\frac{1}{|V|}\sum_{i\in V}F_{p_{i}},$
    where $F_{p_{i}}=-\sum_{s}p_{is}\log p_{is}$ represents the entropy of node $i$’s neighbors, and $p_{is}$ denotes the probability that node $i$’s neighbors belong to subgroup $s$. A higher value indicates a more balanced neighborhood distribution, while a lower value indicates bias toward certain subgroups.
Fairness Evaluation Metrics
We evaluate fairness by using Statistical Parity (SP) (Dwork et al. 2012) and Equality of Opportunity (EO) (Hardt, Price, and Srebro 2016), where lower values indicate better model fairness. Specifically, SP measures the difference in prediction rates between subgroups with different sensitive attributes:
$\Delta SP=|P(\mathbf{\hat{y}}=1|\mathbf{s}=0)-P(\mathbf{\hat{y}}=1|\mathbf{s}=1)|,$ (12)
where $\mathbf{\hat{y}}$ represents the predicted labels and $\mathbf{s}$ indicates the sensitive attribute values.
On the other hand, EO measures the difference in true positive rates across different sensitive attributes:
$\Delta EO=|P(\mathbf{\hat{y}}=1|\mathbf{s}=0,\mathbf{y}=1)-P(\mathbf{\hat{y}}=1|\mathbf{s}=1,\mathbf{y}=1)|,$ (13)
where $\mathbf{y}$ indicates the actual labels.
Appendix B Detailed Implementation Details
The experiments were run on a Linux system with a 36-core CPU, RTX A6000 48GB GPU, and 256GB of RAM. All experiments were conducted by using five different seed settings, and we report the average accuracy. We used GCN as the backbone for both the competitors and DAB-GNN, carefully tuning their hyperparameters via grid search.
For DAB-GNN, we used a 3-layer GCN with a hidden layer of size 16, a gradient penalty (Gulrajani et al. 2017) of 10, 1,000 epochs, the embedding dimensionality of 48, and a weight decay of 1e-5. The optimal values for the remaining hyperparameters are as follows:
  • Learning Rates: 0.001 for AbDisen (all datasets); 0.0003 for SbDisen (all datasets); 0.0005 for PbDisen (NBA) and 0.003 for PbDisen (Recidivism, Credit, Pokec_n, Pokec_z)
  • $\boldsymbol{\alpha}$: 1 (NBA, Recidivism, Pokec_n), 15 (Credit), 100 (Pokec_z)
  • $\boldsymbol{\beta}$: 0.0003 (NBA, Pokec_z), 0.0005 (Recidivism), 0.005 (Pokec_n), 0.0095 (Credit)
  • $\boldsymbol{k}$: 1 (Recidivism), 8 (Credit), 40 (NBA), 100 (Pokec_n, Pokec_z)
For hyperparameters of competitors, we used the best settings found via grid search. We tuned the hyperparameters over the following ranges:
  • FairGNN: $\alpha\in$0.0001, 0.001, 0.01, 0.1, 1} and $\beta\in$1, 10, 20, 30, 40, 50, 60, 70, 80, 90, 100}
[…]
  • EDITS: $u_{1}$in {0.01, 0.03, 0.05, 0.07, 0.1, 0.3, 0.5, 0.7, 1} and $u_{2}$in {0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9, 1}; the proportion was tuned among the values {0.02, 0.29, 0.015}
[…]
  • BIND: we selected the best trade-off value among the available budgets
[…]
(a) NBA
(b) Pokec_n
(c) Pokec_z
Figure I: Visualization of disentangled node embeddings by using t-SNE: attribute (purple), structure (red), and potential (yellow) bias embeddings.
Appendix C Further Experimental Results
Disentangled Embeddings Analysis
Figure I shows the results of visualizing the disentangled embeddings on NBA, Pokec_z, and Pokec_n datasets.
Hyperparameter Analysis
Figure II shows the results of the hyperparameters $\alpha$ and $\beta$, while Figure III shows the results of a hyperparameter $k$. These results are displayed on the following pages.
(a) AUC (b) SP (c) EO
Refer to caption Refer to caption Refer to caption
(i) NBA
Refer to caption Refer to caption Refer to caption
(ii) Recidivism
Refer to caption Refer to caption Refer to caption
(iii) Credit
Refer to caption Refer to caption Refer to caption
(iv) Pokec_n
Refer to caption Refer to caption Refer to caption
(v) Pokec_z
Figure II: The effects of $\alpha$ and $\beta$ on AUC ($\uparrow$), SP ($\downarrow$), and EO ($\downarrow$) across different datasets.
(a) AUC (b) SP (c) EO
Refer to caption Refer to caption Refer to caption
(i) NBA
Refer to caption Refer to caption Refer to caption
(ii) Recidivism
Refer to caption Refer to caption Refer to caption
(iii) Credit
Refer to caption Refer to caption Refer to caption
(iv) Pokec_n
Refer to caption Refer to caption Refer to caption
(v) Pokec_z
Figure III: The effects of $k$ on AUC ($\uparrow$), SP ($\downarrow$), and EO ($\downarrow$) across different datasets.

### v2

@@ 4 Evaluation
We also showed that DAB-GNN requires a reasonable computational cost, with training time increasing approximately linearly as the number of nodes grows. Detailed experimental results are available in the online appendix at https://github.com/Bigdasgit/DAB-GNN.
@@
Datasets. We used 5 real-world graph datasets for our experiments: NBA, Recidivism, Credit, Pokec_n, and Pokec_z, which are all publicly available. Table 1 provides key statistics for these datasets (more details in Section 2).
@@
Evaluation Tasks. Following previous studies (Agarwal, Lakkaraju, and Zitnik 2021; Dai and Wang 2021; Dong et al. 2022; Wang et al. 2022), we assessed the methods by using a node classification task, splitting the nodes into training (50%), validation (25%), and test (25%) sets. We measured accuracy with three metrics: Accuracy (ACC), Area Under the Curve (AUC), and F1-Score. In addition, we evaluate fairness by using Statistical Parity (SP) (Dwork et al. 2012) and Equality of Opportunity (EO) (Hardt, Price, and Srebro 2016), where lower values indicate better model fairness. Detailed equations for calculating SP and EO are provided in the online appendix.
Implementation Details. We used Vanilla GCN as the backbone for both the competitors and DAB-GNN, carefully tuning their hyperparameters via grid search. For DAB-GNN, we used a 3-layer GCN with a hidden layer of size 16, a gradient penalty (Gulrajani et al. 2017) of 10, 1,000 epochs, the embedding dimensionality of 48, and a weight decay of 0.00001. The hyperparameters $k$ for $k$-NN graph construction and $\alpha$ for $\mathcal{L}_{\text{fh}}$ were tuned within the range of {1, 100}, while $\beta$ for $\mathcal{L}_{\text{bco}}$ was tuned within {0.00001, 0.1}. All experiments were conducted by using five different seed settings, and we report the average accuracy. For complete implementation details, please refer to the online appendix.
@@ Results
Metrics (a) Disen. and Amp. Module (b) Debiasing Module DAB-GNN
NBA ACC ($\uparrow$) 70.38 70.89 71.65 69.11 69.37 71.39
AUC ($\uparrow$) 79.51 81.13 81.86 77.68 81.14 80.56
F1 ($\uparrow$) 73.14 74.07 73.36 70.48 72.05 73.51
SP ($\downarrow$) 6.08 5.97 6.52 9.50 6.76 1.12
EO ($\downarrow$) 8.66 8.26 12.25 12.65 8.03 0.80
Recidivism ACC ($\uparrow$) 88.82 91.59 91.25 91.63 91.01 89.99
AUC ($\uparrow$) 92.38 94.13 94.07 94.24 94.04 93.41
F1 ($\uparrow$) 84.20 88.17 87.72 87.73 87.53 86.31
SP ($\downarrow$) 1.13 1.21 1.03 3.30 0.86 0.73
EO ($\downarrow$) 0.77 2.18 1.45 2.26 1.61 0.90
Credit ACC ($\uparrow$) 74.22 75.06 75.75 76.47 74.10 78.19
AUC ($\uparrow$) 72.86 68.45 72.05 64.08 69.66 71.41
F1 ($\uparrow$) 82.84 84.04 84.80 85.93 83.16 87.39
SP ($\downarrow$) 6.89 5.68 4.73 2.59 5.98 0.44
EO ($\downarrow$) 5.67 4.90 3.76 2.30 5.12 0.45
Pokec_n ACC ($\uparrow$) 61.18 67.33 66.50 66.79 67.75 67.18
AUC ($\uparrow$) 67.87 72.89 72.96 73.62 72.82 73.68
F1 ($\uparrow$) 58.22 61.48 63.85 64.54 62.78 62.34
SP ($\downarrow$) 2.03 2.21 2.04 4.05 1.38 0.71
EO ($\downarrow$) 2.57 3.30 2.64 5.40 3.61 1.09
Pokec_z ACC ($\uparrow$) 67.32 68.92 68.82 69.02 67.99 68.56
AUC ($\uparrow$) 73.70 75.14 74.49 75.39 75.05 74.85
F1 ($\uparrow$) 65.89 67.60 67.16 68.30 69.17 67.94
SP ($\downarrow$) 3.55 1.88 4.82 4.59 2.43 0.67
EO ($\downarrow$) 3.70 1.94 3.97 3.61 2.51 0.73
Table 5: The ablation studies for DAB-GNN.
@@
Figure 2: Visualization of disentangled node embeddings by using t-SNE: AbEmb, SbEmb, and PbEmb.
@@
(a) Recidivism (b) Credit
Figure 3: The effects of $\alpha$ and $\beta$ on AUC ($\uparrow$) and EO ($\downarrow$).
@@ 5 Conclusion
In this work, we identified a significant challenge inherent to existing fairness-aware GNN methods: the entanglement of different bias types in the final node embeddings leads to difficulty in their comprehensive debiasing. To address this challenge, we introduced DAB-GNN, a novel GNN framework that disentangles, amplifies, and debiases the attribute, structure, and potential biases within node embeddings. Extensive experiments on five real-world graph datasets show that DAB-GNN outperforms ten state-of-the-art competitors in balancing accuracy and fairness, while validating the effectiveness of our design choices.
Acknowledgments
The work of Sang-Wook Kim was supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2022-0-00352, No.RS-2022-00155586; A High-Performance Big-Hypergraph Mining Platform for Real-World Downstream Tasks). Yeon-Chang Lee’s work was supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant, funded by the Korea government (MSIT) (No. RS-2020-II201336, Artificial Intelligence Graduate School Program (UNIST)).